We have 3 concentric metallic spheres of radius r1, r2 and r3 having charges q1, q2 and q3 on them. The innermost and outermost spheres are earthed. Find the distribution of charge.

Grade:12th pass

1 Answers

Vikas TU
7 years ago
Innermost is earthed hence its potetial becomes zero.
V = 0 = Kq1/r1 = 0
that is q1 = 0
At middle layer
potential is => Kq2/r2 + 0 = V2
At outermost layer,
V3 = kq3/r3 + V => kq3/r3 + kq2/r2 = 0
hence,
q3/r3 = q2/r2
Final charge,
q1 = 0
q2 = q2
and
q3 = q2*r2/r3